A 59-year-old man has been admitted. He has a 90-minute history of severe crushing chest pain. Received nitroglycercin  1 tablet sublingual stat, nitroglycercin patch to upper extremities every six hours, morphine sulfate 2 mg IV. Timpotic eye drops OU bid. Digoxin 0.25 mg po stat.

1. How should the nurse safely administer transdermal or topical medications?

2. List four principles for administering eye instillations.

3. Identify four sites for subcutaneous injections.

4. Identify three sites for Intramuscular injection sites.

5. Explain the purpose for receiving Morphine Sulfate IV.

Conversions:

A. gr 1/2 = ___ mg

B. 2 teaspoons = mL

C. 6 ounces____=mL

50 mL NS with antibiotic IV to infuse in 20 minutes. The drop factor is 20 gtt/mL. Flow rate: ______ gtt/min
Answer #1

1. Nursing measures in safe administration of topical medications.

                * After identification of patient expain the procedure and wash hands and wear glove

                  on dominant hand to prevent spread of micro organisms.

                * Expose only the area where oilment or cream is applied to prevent undue exposure.

                * Clean area with soap and water to remove microorganisms in that area and also

                   remove previously applied medications in that area.

                * Follow the 10 rights of medication administration (Right drug,Right patient,Right route.

                  Right time and frequency,Right documentation,Right history and assessment,Right to

                  refuse,Right drug to drug interaction and evaluation and Right education.) before drug

                  administration.

2. Four principles of administering eye installation

                * Check the opthalmic preparation for name,strength,number of drops and expiry date

                  before instillation.

                * Clean the eyelids by wiping from inner canthus of eye to outer canthus

                * Close the eyelid and avoid squeezing of eyelids after administration of medication.

                * Press the nasolacrymal duct for atleast 30 seconds after instillation of eye medication.

3. Four sites of subcutaneous injection

                * Outer aspect of upper arm.

                * Abdomen near umbilical area.

                * Anterior thigh.(four inch below the top of thigh to four inch above the knee).

                * Upper back of buttocks (behind the hip bone).

4.Three sites of intramuscular injection.

                 * Deltoid site.

                 * Ventrogluteal site.

                 * Vastus lateralis site.

5. Purpose of receiving morphine sulphate IV.

                 * Morphine sulphate is a type of opioid analgesic which used to treat severe pain.

                 * It take only 20 minutes to make maximum effect when giving intravenously.

* Conversions

         A. gr 1/2 = 32.39 mg. ( 1 grain = 64 .79 mg)

         B. 2 Teaspoons = 9.86 ml.( 1 teaspoon = 4.92ml)

         C. 6 ounces = 177.4 ml. ( 1 ounce = 29.57 ml)

* Flow rate

           Formula

                 Flow rate = Volume (ml) / Time (min) * Drop factor (gtts / ml).

                    Here

                        Volume = 50 ml.

                         Time = 20 minutes.

                        Drop factor = 20 gtt /ml.

                Flow rate = 50 / 20 * 20

                                = 2.5 * 20

                                = 50 gtt / min.